Constants Part A The electric field at a distance of 0.154 m from the surface of a solid insulating sphere with radius 0.364 m is 1750 N/C Assuming the spheres charge is uniformly distributed, what is the charge density inside it? You may want to review (Pages 738-742) For related may want to view a Video Tutor Solution of Field of a uniformly charged sphere g tips and strategies, you Submit Incorrect; Try Again; One attempt remaining Part EB Calculate the electric field inside the sphere at a distance of 0.241 m from the center.
Answer #1

Solution)

Part a) Given, Radius of the sphere R = 0.364 m

Distance of the point from the surface of the sphere x = 0.154 m

Distance of the point from center of the sphere r = R + x

= 0.518 m

Electric field , E = 1750 N / C

We know,

E = Kq / r ^2

So, q = Er^ 2 / K

Now, on substituting values we get

q = 5.217 x 10 -8 C

Volume of the sphere V = ( 4/ 3) πR ^3

= 0.2019 m 3

So, Charge density ρ = q / V

= 2.58x10^-7C / m 3 (Ans)

======

Good luck!:)